Evaluate the line integral βˆ«π’–(π‘₯, 𝑦, 𝑧) Γ— ⅆ𝒓 𝐢 , where 𝒖(π‘₯, 𝑦, 𝑧) = (𝑦 2 , π‘₯, 𝑧) and the curve π‘ͺ is described by 𝒛 = 𝑦 = 𝑒 π‘₯ with π‘₯ ∈ [0,1].


1
Expert's answer
2021-11-26T15:12:00-0500

"Since,\\
z=e^x\\
dz=e^xdx\\
And,\\
y=e^x\\
dy=e^xdx\\
Therefore,\\
dr=(dx,dy,dz)=(dx,e^xdx,e^xdx)=(1,e^x,e^x)dx\\
Then,\\
\int u.dr\\
=\int(y^2,x,z).(1,e^x,e^x)dx\\
=\int(e^{2x},x,e^x).(1,e^x,e^x)dx\\
=\int_0^1(e^{2x}+xe^x+e^{2x})dx\\
=\int_0^1(xe^x+2e^{2x})dx\\
=[xe^x-e^x+e^{2x}]_0^1\\
=e^2"